OpenStudy (anonymous):

evaluate the integral: [-1.1] f(x)dx where f(x) = {1/2 for x<0} {x^2 for >= 0}

OpenStudy (anonymous):

\[\large \int_{-1}^0\frac{1}{2}dx+\int_0^1x^2dx\]

OpenStudy (anonymous):

answer for that question is 5/6 i just dont know how to get to that

OpenStudy (anonymous):

\[\int\limits_{-1}^{1}f(x)\] where f(x) = \[\left\{ \frac{ 1 }{ 2} \right\}\] for x<0